WHAT IS THE EXACT VALUE OF THE FOLLOWING?
1) SIN 9PI/2
2) COSECANT 7PI

Answer by Alan3354(30993)

WHAT IS THE EXACT VALUE OF THE FOLLOWING?
1) SIN 9PI/2
2) COSECANT 7PI
--------------
This one is old, your need for it might have passed, but...
1) sin(9*PI/2)
2*PI is a complete revolution, so subtract that until the value is less than 2PI.
sin(9PI/2) + sin([9PI/2 - 2PI -2PI]) = sin(PI/2)
= 1.
If it's easier to follow use degrees.
9PI/2 = 9*180/2 deg = 810 degs
810 - 360 = 450
450 - 360 = 90
sin(90) = 1
-----------------
csc(7PI)
Again, subtract 2PI as many times as possible while remaining positive.
csc(7PI) = csc(PI)
csc(PI) = 1/[sin(PI)] = 1/0
it's infinity, or undefined.